# Geodesic Synthetic Control Methods for Random Objects and Functional Data

### v1

@@ A.2. Space of compositional data with the arc-length distance
Next we show (A.3). Specifically, we establihs (i) $\mathbb{W}_{t}^{*}\subset\overline{\mathbb{W}}_{t}^{*}$ and (ii) $\mathbb{W}_{t}^{*}\supset\overline{\mathbb{W}}_{t}^{*}$.
Proof of (i): Fix an arbitrary $\bm{w}=(w_{2},\dots,w_{J+1})^{\prime}\in\mathbb{W}_{t}^{*}$. By definition, $g_{t}^{(\bm{w})}(U_{2:J+1})$ is a Fréchet mean of the distribution a random element $X$ such that $\mathrm{P}(X=g_{t}(U_{j}))=w_{j}$, $j=2,\dots,J+1$. Further, the condition $d(g_{t}(U_{1}),g_{t}^{(\bm{w})}(U_{2:J+1}))=0$ yields $g_{t}(U_{1})=g_{t}^{(\bm{w})}(U_{2:J+1})$. Applying Corollary 3 in Le and Barden, (2014), we then have $\sum_{j=2}^{J+1}w_{j}\mathrm{Log}_{g_{t}(U_{1})}(g_{t}(U_{j}))=0$. This implies $\bm{w}\in\overline{\mathbb{W}}_{t}^{*}$ and hence $\mathbb{W}_{t}^{*}\subset\overline{\mathbb{W}}_{t}^{*}$ holds.
Proof of (ii): Fix an arbitrary $\bm{w}=(w_{2},\dots,w_{J+1})^{\prime}\in\overline{\mathbb{W}}_{t}^{*}$. From the definition of $\overline{\mathbb{W}}_{t}^{*}$ and the uniqueness of the Fréchet mean, Theorem 1 in Buss and Fillmore, (2001) yields $g_{t}(U_{1})=g_{t}^{(\bm{w})}(U_{2:J+1})$. This implies $\bm{w}\in\mathbb{W}_{t}^{*}$ and hence $\mathbb{W}_{t}^{*}\supset\overline{\mathbb{W}}_{t}^{*}$.
A.3. Space of SPD matrices
First, we consider the space of SPD matrices with the Frobenius metric. Standard arguments lead to (A.5) and (A.2). Hence Assumption 3.1 holds.
@@ A.3. Space of SPD matrices
Note that the space $(\mathcal{L}^{+}_{m},d_{\mathrm{LC}})$ is a geodesic metric space (Lin,, 2019) and the geodesic with respect to $d_{\mathrm{LC}}$ between $L_{1}$ and $L_{2}$ is
@@
$\log(\mathbb{D}(g_{t}^{(\bm{w})}(U_{2:J+1})))$ $=\log(\mathbb{D}(\mu_{t}))+\alpha_{t}\left\{\sum_{j=2}^{J+1}w_{j}\log(\mathbb{D}(U_{j}))-\log(\mathbb{D}(\mu_{t}))\right\}=\log(\mathbb{D}(g_{t}(U_{2:J+1}^{(\bm{w})}))).$
@@
$d_{\mathrm{LC}}^{2}(Y_{1,t}(N),Y_{2:J+1,t}^{(\bm{w})})$ $=\|\lfloor g_{t}(U_{1})\rfloor-\lfloor g_{t}(U_{2:J+1}^{(\bm{w})})\rfloor\|^{2}_{F}+\|\log(\mathbb{D}(g_{t}(U_{1})))-\log(\mathbb{D}(g_{t}(U_{2:J+1}^{(\bm{w})})))\|^{2}_{F}$
@@
A.4. Space of one-dimensional probability distributions with the 2-Wasserstein distance
@@
A.5. Space of functional data with the $L^{2}$ metric
@@
[…]
@@
Appendix B Proofs
B.1. Proof of Theorem 3.1
@@ Appendix B Proofs
$\overline{\bm{w}}\in\mathop{\rm arg~{}min}\limits_{\bm{w}\in\Delta^{J-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}d^{2}(g_{t}(U_{1}),g_{t}^{(\bm{w})}(U_{2:J+1})).$
@@
B.3. Proof of Theorem 6.1
Note that under Assumption 6.1, we have
@@
To show (6.4) under Assumption 6.2, observe that
$\overline{\bm{w}}\in\mathop{\rm arg~{}min}\limits_{\bm{w}\in\Delta^{J-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}d^{2}(g(U_{1},V_{t}),g^{(\bm{w})}(U_{2:J+1},V_{t})).$
@@
To show (6.4) under Assumption 6.3, observe that
$\overline{\bm{\lambda}}\in\mathop{\rm arg~{}min}\limits_{\bm{\lambda}\in\Delta^{T_{0}-1}}\sum_{j=2}^{J+1}d^{2}(g^{(\oplus)}(U_{j},V_{T_{0}+1:T}),g^{(\bm{\lambda})}(U_{j},V_{1:T_{0}})).$
Then Assumption 6.3 yields $\overline{\bm{\lambda}}\in\bigcap_{j=1}^{J+1}\Lambda_{j}^{*}$ and hence

### v2

@@ 7 Data Availability Statement
Next we show (S7). Specifically, we establihs (i) $\mathbb{W}_{t}^{*}\subset\overline{\mathbb{W}}_{t}^{*}$ and (ii) $\mathbb{W}_{t}^{*}\supset\overline{\mathbb{W}}_{t}^{*}$.
Proof of (i): Fix an arbitrary $\bm{w}=(w_{2},\dots,w_{J+1})^{\prime}\in\mathbb{W}_{t}^{*}$. By definition, $g_{t}^{(\bm{w})}(U_{2:J+1})$ is a Fréchet mean of the distribution a random element $X$ such that $\Pr(X=g_{t}(U_{j}))=w_{j}$, $j=2,\dots,J+1$. Further, the condition $d(g_{t}(U_{1}),g_{t}^{(\bm{w})}(U_{2:J+1}))=0$ yields $g_{t}(U_{1})=g_{t}^{(\bm{w})}(U_{2:J+1})$. Applying Corollary 3 in Le and Barden (2014), we then have $\sum_{j=2}^{J+1}w_{j}\mathrm{Log}_{g_{t}(U_{1})}(g_{t}(U_{j}))=0$. This implies $\bm{w}\in\overline{\mathbb{W}}_{t}^{*}$ and hence $\mathbb{W}_{t}^{*}\subset\overline{\mathbb{W}}_{t}^{*}$ holds.
Proof of (ii): Fix an arbitrary $\bm{w}=(w_{2},\dots,w_{J+1})^{\prime}\in\overline{\mathbb{W}}_{t}^{*}$. From the definition of $\overline{\mathbb{W}}_{t}^{*}$ and the uniqueness of the Fréchet mean, Theorem 1 in Buss and Fillmore (2001) yields $g_{t}(U_{1})=g_{t}^{(\bm{w})}(U_{2:J+1})$. This implies $\bm{w}\in\mathbb{W}_{t}^{*}$ and hence $\mathbb{W}_{t}^{*}\supset\overline{\mathbb{W}}_{t}^{*}$.
S.4.3 Space of SPD matrices
First, we consider the space of SPD matrices with the Frobenius metric. Standard arguments lead to (S.4.5) and (S6). Hence Assumption 3.1 holds.
@@
Note that the space $(\mathcal{L}^{+}_{m},d_{\mathrm{LC}})$ is a geodesic metric space (Lin, 2019) and the geodesic with respect to $d_{\mathrm{LC}}$ between $L_{1}$ and $L_{2}$ is
@@
$\log(\mathbb{D}(g_{t}^{(\bm{w})}(U_{2:J+1})))$ $=\log(\mathbb{D}(\mu_{t}))+\alpha_{t}\left\{\sum_{j=2}^{J+1}w_{j}\log(\mathbb{D}(U_{j}))-\log(\mathbb{D}(\mu_{t}))\right\}$
$=\log(\mathbb{D}(g_{t}(U_{2:J+1}^{(\bm{w})}))).$
@@
$d_{\mathrm{LC}}^{2}(Y_{1,t}(N),Y_{2:J+1,t}^{(\bm{w})})$
$=\|\lfloor g_{t}(U_{1})\rfloor-\lfloor g_{t}(U_{2:J+1}^{(\bm{w})})\rfloor\|^{2}_{F}+\|\log(\mathbb{D}(g_{t}(U_{1})))-\log(\mathbb{D}(g_{t}(U_{2:J+1}^{(\bm{w})})))\|^{2}_{F}$
@@
S.4.4 Space of univariate probability distributions with the 2-Wasserstein distance
@@
S.4.5 Space of functional data with the $L^{2}$ metric
@@
$d_{L^{2}}^{2}(Y_{1,t}(N),Y_{2:J+1,t}^{(\bm{w})})$ $=d_{L^{2}}(g_{t}(U_{1}),g_{t}(U_{2:J+1}^{(\bm{w})}))=\int_{\mathcal{T}}\left\{(g_{t}(U_{1}))(s)-(g_{t}(U_{2:J+1}^{(\bm{w})}))(s)\right\}^{2}ds$
$=\alpha_{t}^{2}\int_{\mathcal{T}}\left\{U_{1}(s)-U_{2:J+1}^{(\bm{w})}(s)\right\}^{2}ds=\alpha_{t}^{2}d_{L^{2}}(U_{1},U_{2:J+1}^{(\bm{w})}),$
@@
S.5 Proofs
S.5.1 Proof of Theorem 3.1
@@
$\overline{\bm{w}}\in\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}d^{2}(g_{t}(U_{1}),g_{t}^{(\bm{w})}(U_{2:J+1})).$
@@
S.5.3 Proof of Theorem S1
Note that under Assumption S2, we have
@@
To show (S4) under Assumption S3, observe that
$\overline{\bm{w}}\in\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}d^{2}(g(U_{1},V_{t}),g^{(\bm{w})}(U_{2:J+1},V_{t})).$
@@
To show (S4) under Assumption S4, observe that
$\overline{\bm{\lambda}}\in\operatorname*{arg\,min}_{\bm{\lambda}\in\Delta^{T_{0}-1}}\sum_{j=2}^{J+1}d^{2}(g^{(\oplus)}(U_{j},V_{T_{0}+1:T}),g^{(\bm{\lambda})}(U_{j},V_{1:T_{0}})).$
Then Assumption S4 yields $\overline{\bm{\lambda}}\in\bigcap_{j=1}^{J+1}\Lambda_{j}^{*}$ and hence